1. Determine the first order partial derivative of the following functions:

a. z = In (x + t^2)

b. F(x,y,z) = xy^2 e^-xz


2. Clairaut's Theorwm holds that Uxy = Uyx , show that the following equations obey Clairaut's Theorem:


a. U = In( x+2y)

b. U = e^xy sin y


3. Laplaces equation holds that Uxx + Uyy= 0, verify that the second derivative of the following equations are Laplace's equations:


a. U = In√x^2 + y^2

b. U = x^2 - y^2




1
Expert's answer
2022-01-10T14:36:41-0500

1.

a.

"\dfrac{\partial z}{\partial x}=\dfrac{1}{x+t^2}"

"\dfrac{\partial z}{\partial t}=\dfrac{2t}{x+t^2}"

b.


"\dfrac{\partial F}{\partial x}=y^2e^{-xz}-xy^2ze^{-xz}""\dfrac{\partial F}{\partial y}=2ye^{-xz}""\dfrac{\partial F}{\partial z}=-x^2y^2e^{-xz}"

2.

a.


"\dfrac{\partial U}{\partial x}=\dfrac{1}{x+2y}"

"\dfrac{\partial U}{\partial y\partial x}=-\dfrac{2}{(x+2y)^2}"


"\dfrac{\partial U}{\partial y}=\dfrac{2}{x+2y}"

"\dfrac{\partial U}{\partial x\partial y}=-\dfrac{2}{(x+2y)^2}"

"\dfrac{\partial U}{\partial y\partial x}=-\dfrac{2}{(x+2y)^2}=\dfrac{\partial U}{\partial x\partial y}, True"

b.


"\dfrac{\partial U}{\partial x}=ye^{xy}\sin y"

"\dfrac{\partial U}{\partial y\partial x}=e^{xy}\sin y+xye^{xy}\sin y+ye^{xy}\cos y"


"\dfrac{\partial U}{\partial y}=xe^{xy}\sin y+e^{xy}\cos y"

"\dfrac{\partial U}{\partial x\partial y}=e^{xy}\sin y+xye^{xy}\sin y+ye^{xy}\cos y"

"\dfrac{\partial U}{\partial y\partial x}=e^{xy}\sin y+xye^{xy}\sin y+ye^{xy}\cos y"




"=\dfrac{\partial U}{\partial x\partial y}, True"

3.

a.

"U = \ln\sqrt{x^2 + y^2}=\dfrac{1}{2}\ln (x^2+y^2)"

"U_x=\dfrac{1}{2}(\dfrac{1}{x^2+y^2})(2x)=\dfrac{x}{x^2+y^2}"

"U_{xx}=\dfrac{x^2+y^2-2x^2}{(x^2+y^2)^2}=\dfrac{y^2-x^2}{(x^2+y^2)^2}"

"U_y=\dfrac{1}{2}(\dfrac{1}{x^2+y^2})(2x)=\dfrac{y}{x^2+y^2}"

"U_{yy}=\dfrac{x^2+y^2-2y^2}{(x^2+y^2)^2}=\dfrac{x^2-y^2}{(x^2+y^2)^2}"

"U_{xx}+U_{yy}=\dfrac{y^2-x^2}{(x^2+y^2)^2}+\dfrac{x^2-y^2}{(x^2+y^2)^2}=0, True"

b.

"U_x=2x"

"U_{xx}=2"

"U_y=-2y"

"U_{yy}=-2"

"U_{xx}+U_{yy}=2-2=0, True"